OpenStudy (anonymous):

Solve the following system of equations. Express your answer as an ordered pair in the format (a,b), with no spaces between the numbers or symbols. 2x+7y=-1 4x-3y=-19

OpenStudy (anonymous):

Solve the system by whatever method you prefer (substitution, elimination, matricies, etc.) and then express your answer as (x,y), using the solution for x and y you found previously. Does solving the system give you trouble or the format of the answer?

OpenStudy (anonymous):

solving the system confuses me

OpenStudy (anonymous):

I would probably use elimination to solve this one. \[2x+7y=-1\]\[4x-3y=-19\]Multiply the first equation by -2 and add it to the 2nd equation. \[-2(2x)+4x-2(7y)-3y=-2(-1)-19\]\[-17y=-17\]Now solve for y. Then plug in your answer for y into either the 1st or 2nd equation to find x.